# Naoyuki Kamiyama

### v1

@@ Abstract
The stable roommates problem is a non-bipartite version of the well-known stable matching problem. Teo and Sethuraman proved that, for each instance of the stable roommates problem in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. The aim of this paper is to extend the result of Teo and Sethuraman to the stable roommates problem with ties. More concretely, we prove that, for each instance of the stable roommates problem with ties in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance.
@@ 1 Introduction
The stable roommates problem is a non-bipartite version of the stable matching problem in a bipartite graph [7]. In contrast to the stable matching problem, it is known that there exists an instance of the stable roommates problem where a stable matching does not exist [7, Example 3]. Thus, the problem of checking the existence of a stable matching in a given instance of the stable roommates problem is one of the most important problems in the study of the stable roommates problem. For this problem, Irving [11] proposed a polynomial-time algorithm. This algorithm is combinatorial, i.e., this does not need solving linear programs. On the other hand, in the study of the stable matching problem, linear programming approaches have been actively studied (see, e.g., [4, 13, 17]). Thus, it is natural to investigate a linear programming approach to the stable roommates problem. In this direction, Teo and Sethuraman [15] proposed the following linear programming approach. They proved that, for each instance of the stable roommates problem in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. Since a linear program such that the separation problem can be solved in polynomial time can be solved in polynomial time [8], the result of Teo and Sethuraman [15] can be regarded as another proof of the polynomial-time solvability of the problem of checking the existence of a stable matching in a given instance of the stable roommates problem in the complete graphs. (See [14, 16] for related work of [15] and [1, 2] for another linear programming approach to the stable roommates problem.)
In this paper, we extend the linear programming approach of Teo and Sethuraman [15] to the stable roommates problem with ties. In particular, we focus on super-stability in the stable roommates problem with ties (see Section 2 for its formal definition). Irving and Manlove [12] proposed a combinatorial polynomial-time algorithm for the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties (see also [5, 6]). In this paper, we prove that, for each instance of the stable roommates problem with ties in the complete graphs, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance. Furthermore, we prove that the separation problem for the above linear inequality system can be solved in polynomial time, and a super-stable matching can be constructed from a feasible solution to the system in polynomial time. Thus, the result of this paper gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in the complete graphs.
@@ 2 Preliminaries
Let $\mathbb{R}_{+}$ denote the set of non-negative real numbers. For each finite set $U$, each vector $x\in\mathbb{R}_{+}^{U}$, and each subset $W\subseteq U$, we define $x(W):=\sum_{u\in W}x(u)$. For each positive integer $z$, we define $[z]:=\{1,2,\dots,z\}$.
In this paper, we are given the complete graph $G=(V,E)$ on a finite vertex set $V$ such that $|V|$ is an even positive integer. Notice that $E=\{e\subseteq V\mid|e|=2\}$. For each subset $F\subseteq E$ and each vertex $v\in V$, we define $F(v)\coloneqq\{e\in F\mid v\in e\}$. For each vertex $v\in V$, we are given a transitive binary relation $\succsim_{v}$ on $E(v)$ such that, for each pair of edges $e,f\in E(v)$, at least one of $e\succsim_{v}f$, $f\succsim_{v}e$ holds. For each vertex $v\in V$ and each pair of edges $e,f\in E(v)$, if $e\succsim_{v}f$ and $f\not\succsim_{v}e$ (resp. $e\succsim_{v}f$ and $f\succsim_{v}e$), then we write $e\succ_{v}f$ (resp. $e\sim_{v}f$). Intuitively speaking, if $e\succ_{v}f$, then $v$ prefers $e$ to $f$. If $e\sim_{v}f$, then $v$ is indifferent between $e$ and $f$.
@@
Let $\mu$ be a matching in $G$. Then for each edge $e\in E\setminus\mu$ and each vertex $v\in e$, we say that $e$ weakly blocks $\mu$ on $v$ if $e\succsim_{v}\mu(v)$. Furthermore, for each edge $e\in E\setminus\mu$, we say that $e$ weakly blocks $\mu$ if $e$ weakly blocks $\mu$ on $v$ for every vertex $v\in e$.
@@
If the preferences do not contain ties, then Theorem 1 coincides with [15, Theorem 4]. Since we can prove that the separation problem for ${\bf P}$ can be solved in polynomial time in a similar way as [15] (see Section 4), Theorem 1 gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in the complete graphs. Furthermore, the proof of Theorem 1 implies that if ${\bf P}\neq\emptyset$, then a super-stable matching can be constructed from an element in ${\bf P}$ in polynomial time.
@@ 3 Proof of Theorem 1
For each vertex $v\in V$, an edge $e\in S(v)$ satisfying
@@
For every vertex $v\in V$ such that $x(E[\succsim_{v}d(v)])>1/2$, $m(m(v))=v$.
@@
Let $v$ be a vertex in $V$. Define $w\coloneqq m(v)$. Then $d(v)=\{v,w\}$. Assume that $m(w)\neq v$. Thus, since $d(w)\neq d(v)$, we have at least one of $x(E[\succsim_{w}d(v)])<1/2$, $x(E[\succ_{w}d(v)])\geq 1/2$.
@@
Define the subset $L\subseteq E$ as follows. For each pair of distinct vertices $v,w\in V$, $\{v,w\}\in L$ if and only if $m(v)=w$. (Notice that even if $m(v)=w$ and $m(w)=v$ for some pair of distinct vertices $v,w\in V$, $L$ contains a single edge between $v$ and $w$.) Then $\bigcup_{e\in L}e=V$. By considering the directed graph such that its vertex set is $V$ and its arc set contains an arc from $v$ to $m(v)$ for each vertex $v\in V$, we can prove that there exist a set $\mathcal{C}$ of vertex-disjoint cycles in $G$ and a set $M$ of pairwise disjoint edges in $L$ satisfying the following conditions.
@@
Let $\mu$ be a matching in $G$ such that $\mu\subseteq L$, and let $e$ be an edge in $E\setminus L$. Assume that $e$ weakly blocks $\mu(v)$. Then for every vertex $v\in e$, $d(v)\succ_{v}e\succsim_{v}d^{-}(v)$.
@@
Let $v$ be a vertex in $e$. Since $\mu\subseteq L$, $\mu(v)\in\{d(v),d^{-}(v)\}$. If $\mu(v)=d(v)$, then (9) implies that $x(E[\succ_{v}\mu(v)])<1/2$. Assume that $\mu(v)\neq d(v)$. Then $\mu(v)=d^{-}(v)$ and $d(v)\neq d^{-}(v)$. Since $d(v)\neq d^{-}(v)$, we have $m(m(p(v)))\neq p(v)$, i.e., $p(v)\in V\setminus V^{\ast}$. This and Lemma 7 imply that $v\in V\setminus V^{\ast}$. Thus, Lemma 9 implies that $1/2=x(E[\succ_{v}d^{-}(v)])=x(E[\succ_{v}\mu(v)])$. ∎
@@ 4 Polynomial-Time Solvability
Let $x$ be a vector in $\mathbb{R}_{+}^{E}$. It is not difficult to see that we can determine whether $x$ satisfies (1) and (2) in polynomial time. Assume that $x$ satisfies (1) and (2). In what follows, under this assumption, we consider the problem of determining whether $x$ satisfies (3).
@@
$\begin{split}1>{\sf cost}(W)&=\sum_{i=1}^{k}\big{(}1-x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})-x(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\big{)}\\ &=k-2\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\}).\end{split}$
@@
$\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>(k-1)/2=\lfloor k/2\rfloor.$
@@
$\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>\lfloor k/2\rfloor,$ (12)
@@
$\begin{split}{\sf cost}(W)&=k-2\sum_{i=1}^{k}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})\\ &<k-2\lfloor k/2\rfloor=k-2((k-1)/2)=1.\end{split}$

### v2

@@ Abstract
The stable roommates problem is a non-bipartite version of the well-known stable matching problem. Teo and Sethuraman proved that, for each instance of the stable roommates problem in a complete graph, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. The aim of this paper is to extend the result of Teo and Sethuraman to the stable roommates problem with ties. More concretely, we prove that, for each instance of the stable roommates problem with ties in a complete graph, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance.
@@ 1 Introduction
The stable roommates problem is a non-bipartite version of the stable matching problem in a bipartite graph [7]. In contrast to the stable matching problem, it is known that there exists an instance of the stable roommates problem where a stable matching does not exist [7, Example 3]. Thus, the problem of checking the existence of a stable matching in a given instance of the stable roommates problem is one of the most important problems in the study of the stable roommates problem. For this problem, Irving [11] proposed a polynomial-time algorithm. This algorithm is combinatorial, i.e., this does not need solving linear programs. On the other hand, in the study of the stable matching problem, linear programming approaches have been actively studied (see, e.g., [4, 13, 17]). Thus, it is natural to investigate a linear programming approach to the stable roommates problem. In this direction, Teo and Sethuraman [15] proposed the following linear programming approach. They proved that, for each instance of the stable roommates problem in a complete graph, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a stable matching in the given instance. Since a linear program such that the separation problem can be solved in polynomial time can be solved in polynomial time [8], the result of Teo and Sethuraman [15] can be regarded as another proof of the polynomial-time solvability of the problem of checking the existence of a stable matching in a given instance of the stable roommates problem in a complete graph. (See [14, 16] for related work of [15] and [1, 2] for another linear programming approach to the stable roommates problem.)
In this paper, we extend the linear programming approach of Teo and Sethuraman [15] to the stable roommates problem with ties. In particular, we focus on super-stability in the stable roommates problem with ties (see Section 2 for its formal definition). Irving and Manlove [12] proposed a combinatorial polynomial-time algorithm for the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties (see also [5, 6]). In this paper, we prove that, for each instance of the stable roommates problem with ties in a complete graph, there exists a linear inequality system such that there exists a feasible solution to this system if and only if there exists a super-stable matching in the given instance. Furthermore, we prove that the separation problem for the above linear inequality system can be solved in polynomial time, and a super-stable matching can be constructed from a feasible solution to the system in polynomial time. Thus, the result of this paper gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in a complete graph.
@@ 2 Preliminaries
Let $\mathbb{R}_{+}$, $\mathbb{Q}_{+}$ denote the sets of non-negative real numbers and non-negative rational numbers, respectively. For each positive integer $z$, we define $[z]:=\{1,2,\dots,z\}$. For each finite set $U$, each vector $x\in\mathbb{R}_{+}^{U}$, and each subset $W\subseteq U$, we define $x(W):=\sum_{u\in W}x(u)$.
In this paper, we are given the complete graph $G=(V,E)$ on a finite vertex set $V$ such that $|V|$ is an even positive integer. Notice that $E=\{e\subseteq V\mid|e|=2\}$. For each subset $F\subseteq E$ and each vertex $v\in V$, we define $F(v)\coloneqq\{e\in F\mid v\in e\}$. For each vertex $v\in V$, we are given a transitive binary relation $\succsim_{v}$ on $E(v)$ such that, for every pair of edges $e,f\in E(v)$, at least one of $e\succsim_{v}f$, $f\succsim_{v}e$ holds. For each vertex $v\in V$ and each pair of edges $e,f\in E(v)$, if $e\succsim_{v}f$ and $f\not\succsim_{v}e$ (resp. $e\succsim_{v}f$ and $f\succsim_{v}e$), then we write $e\succ_{v}f$ (resp. $e\sim_{v}f$). Intuitively speaking, if $e\succ_{v}f$, then $v$ prefers $e$ to $f$. If $e\sim_{v}f$, then $v$ is indifferent between $e$ and $f$.
@@
Let $\mu$ be a matching in $G$. Then for each edge $e\in E\setminus\mu$, we say that $e$ weakly blocks $\mu$ if $e\succsim_{v}\mu(v)$ for every vertex $v\in e$.
@@
If the preferences do not contain ties, then Theorem 1 coincides with [15, Theorem 4]. Since we can prove that the separation problem for ${\bf P}$ can be solved in polynomial time in a similar way as [15] (see Section 4), Theorem 1 gives another proof of the polynomial-time solvability of the problem of checking the existence of a super-stable matching in a given instance of the stable roommates problem with ties in a complete graph. Furthermore, the proof of Theorem 1 implies that if ${\bf P}\neq\emptyset$, then a super-stable matching can be constructed from an element in ${\bf P}$ in polynomial time.
@@ 3 Proof of Theorem 1
For every vertex $v\in V$, an edge $e\in S(v)$ satisfying
@@
Let $v$ be a vertex in $V$ such that $x(E[\succsim_{v}d(v)])>1/2$. Then $m(m(v))=v$.
@@
Define $w\coloneqq m(v)$. Then $d(v)=\{v,w\}$. Assume that $m(w)\neq v$. Since $d(w)\neq d(v)$, we have at least one of $x(E[\succsim_{w}d(v)])<1/2$, $x(E[\succ_{w}d(v)])\geq 1/2$.
@@
Define the subset $L\subseteq E$ as follows. For each pair of distinct vertices $v,w\in V$, $\{v,w\}\in L$ if and only if at least one of $m(v)=w$, $m(w)=v$ holds. Then $\bigcup_{e\in L}e=V$. By considering the directed graph such that its vertex set is $V$ and its arc set contains an arc from $v$ to $m(v)$ for each vertex $v\in V$, we can prove that there exist a set $\mathcal{C}$ of vertex-disjoint cycles in $G$ and a set $M$ of pairwise disjoint edges in $L$ satisfying the following conditions.
@@
Let $\mu$ be a matching in $G$ such that $\mu\subseteq L$, and let $e$ be an edge in $E\setminus L$. Assume that $e$ weakly blocks $\mu$. Then for every vertex $v\in e$, $d(v)\succ_{v}e\succsim_{v}d^{-}(v)$.
@@
Let $v$ be a vertex in $e$. Since $\mu\subseteq L$, $\mu(v)\in\{d(v),d^{-}(v)\}$. If $\mu(v)=d(v)$, then (9) implies that $x(E[\succ_{v}\mu(v)])<1/2$. Assume that $\mu(v)\neq d(v)$. Then $\mu(v)=d^{-}(v)$ and $d(v)\neq d^{-}(v)$. Since $d(v)\neq d^{-}(v)$, we have $v\in V\setminus V^{\ast}$. Thus, Lemma 9 implies that $1/2=x(E[\succ_{v}d^{-}(v)])=x(E[\succ_{v}\mu(v)])$. ∎
@@ 4 Polynomial-Time Solvability
Let $x$ be a vector in $\mathbb{Q}_{+}^{E}$. It is not difficult to see that we can determine whether $x$ satisfies (1) and (2) in polynomial time. Assume that $x$ satisfies (1) and (2). In what follows, under this assumption, we consider the problem of determining whether $x$ satisfies (3).
@@
$\begin{split}1>{\sf cost}(W)&=\sum_{i\in[k]}\big{(}1-x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})-x(E[e_{i+1}\succsim_{v_{i+1}}]\setminus\{e_{i+2}\})\big{)}\\ &=k-2\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\}).\end{split}$
@@
$\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>(k-1)/2=\lfloor k/2\rfloor.$
@@
$\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})>\lfloor k/2\rfloor,$ (12)
@@
$\begin{split}{\sf cost}(W)&=k-2\sum_{i\in[k]}x(E[e_{i}\succsim_{v_{i}}]\setminus\{e_{i+1}\})\\ &<k-2\lfloor k/2\rfloor=k-2((k-1)/2)=1.\end{split}$